graph each equation
x+y=6
x-y=5
x-y=-3
2x+y=2
x+y=-3
x-2y=6

Answer by ewatrrr(24785)

 
Hi,
graphing equations:
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
x+y=6 Pt(0,6) y = -x +6 m = -1 < 0 line slants to the left
x-y=5 Pt(0,-5)y = x-5
x-y=-3 Pt(0,3) y = x+3
2x+y=2 Pt(0,2) y = -2x+2 m = -2 < 0 line slants to the left
x+y=-3 Pt(0,-3) y = -x-3 m = -1 < 0 line slants to the left
x-2y=6 Pt(0,-3) y = (1/2)x- 3